OpenStudy (anonymous):

Solve for x: −9−(9x−6)=3(4x+6) A. 1 B. -1 C. 3 D. -3

OpenStudy (anonymous):

-9-9x + 6 = 3*4x + 3*6 -9 + 6 - 9x = 12x + 18 -3 - 9x = 12x + 18 12x + 9x = -3-18 21x = -21 x = -1